40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14051     Entry time: Wed Jul 11 15:57:00 2018
Author: aaron 
Type: Update 
Category: OMC 
Subject: Reviving OMC electronics 
Gautam showed me the electronics racks for the OMC PZTs and DAQ. I'm in the process of chasing down what channels we need, and confirming that we'll be able to plug the old antialiasing/imaging boards into the current DAC/ADC boards. I found what I think was Rob Ward's simlink model for the omc, located at
 
/cvs/cds/caltech/cds/rward-advLigo/src/epics/simLink/omc.mdl
 
Channels in this model:
  • 27 or 29 total ADC channels are used (depending whether we keep 2 spare adc chans)
    • 4 each go to ASC_QPD1/2 (8 chans total)
    • 5 go to TRANS_PD1, TRANS_PD2, REFL_PD, TRANS_PD1_UF, TRANS_PD2_UF. These PD are used for ASC and LSC.
    • 2 go to the LSC, one each for DVMDC, DVMAC, X3DC, and X4DC
    • 12 go to the ASC_PZT
    • 2 go to the SPARE_ADC (not sure what this is)
    • I think these channels are (or were at some point) defined in memory by /cvs/cds/caltech/chans/ipc/G1.ipc
      • I found this from elog 2860; it mentions that these should eventually be migrated over to a file C1.ipc, but I don't see any OMC channels in that file or any of the 'old' C1.ipc files, so I suppose it never happened or they were removed later
    • During this vent, we won't have ASC, so
  • 10 or 14 DAC channels are used (depending whether we keep 4 spare dac chans)
    • 2 from the LSC, one for CLK_OUT and one for "LSC"
    • 8 from ASC, including P1A, P1B, P2A, P2B, P1OSC, Y1OSC, P2OSC, Y2OSC
    • I think these channels are (or were at some point) saved to frames due to /cvs/cds/caltech/chans/daq/C1OMC.ini, which I found from elog 2073
    • At some point, the 33MHz mod depth was controlled by one of the spare OMC chans, C1:OMC-SPARE_DAC_CH_15. See elog 2126. I assume this is no longer the case, since c1omc is defunct.
    • Durnig this vent, we won't have ASC and don't need to CLK_OUT the LSC, so we may just need one DAC channel

As of at least Nov 2009, the .par file for the OMC was located at /cvs/cds/gds/param/tpchn_C2 (see elog 2316)

 
Electronics inventory:
  • Kepco HV supply, "OMC-L-PZT", labels indicate it goes to 250V, needs to be tested  ("TESTED OK 2014OCT12")
  • Tip/Tilt Piezo Driver, LIGO D060287
  • HV Piezo Driver, LIGO D060283
  • QPD Whitening Board, D060214
  • LIGO D050374/D050387
  • LIGO D050368/D050373

Need to check:

  • Can the ADC/DAC adapter boards (eg D0902006) drive whatever ~10V control signal we need across ~10m of SCSI cable?
  •  
ELOG V3.1.3-